# A Few Large Shifts: Layer-Inconsistency Based Minimal Overhead Adversarial Example Detection

### v4

@@ Abstract
[…]
@@ 1 Introduction
Existing defensive strategies generally fall into three categories: adversarial training, input purification, and AE detection [13, 1]. Adversarial training, re-trains or fine-tunes a model using adversarially perturbed samples [34, 20], despite its robustness [10, 36], it incurs high computational costs and may compromise clean-data accuracy. Input purification methods [24, 30] attempt to remove adversarial noise through preprocessing (e.g., denoising), yet often fail against adaptive attacks [7].
@@
A more flexible alternative is AE detection, which avoids retraining or significantly modifying the model by rejecting suspicious inputs at test time [35, 40]. Adversarial detection methods can be broadly categorized into two types: those that compare inputs to a reference set or distribution of known patterns (e.g., Deep k-Nearest Neighbors, DkNN [27]; Latent Neighborhood Graph, LNG [1]), and those that analyze invariants in the learned representations [15, 6]. Reference-based detectors often require storing adversarial examples or constructing complex neighbor graphs, which can be computationally intensive and prone to overfitting specific attack types. To address these limitations, recent work [39] proposed detecting consistency between augmented inputs using pre-trained Self-Supervised Learning (SSL) models. Despite avoiding both the need for complex data structures and reliance on adversarial examples, it introduces several practical challenges. Notably, it incurs significant overhead due to the use of large pre-trained SSL models–often significantly larger (e.g., $\times2$ larger params) than the target classifier itself–and assumes the availability of high-quality SSL models for each domain, which may not always be feasible. Furthermore, their method relies solely on SSL models pre-trained on the same datasets used for evaluation.
@@
To eliminate the overhead of external models, complex data structures, or heavy augmentations, we develop a self-contained, layer-wise detector that scrutinizes the network’s own representations. This raises a central question: How do adversarial perturbations propagate through a deep neural network? Prior work shows that such perturbations often leave early-layer features largely intact while inducing sharp deviations in deeper layers [15]. Rather than contrasting these deviations with external reference sets or heavy augmented features, we ask our core research question: Do manifold-induced feature jumps themselves expose adversarial inputs? We formalize this intuition in the A Few Large Shifts Assumption, which posits that adversarial perturbations push representations onto fragile submanifolds at only a few critical layers, producing large, localized shifts. As illustrated in Figure 1, these sparse disruptions leave a detectable footprint across successive layers. Guided by this assumption, we introduce a lightweight detection framework grounded in layer-wise inconsistency. Two complementary probes—Recovery Testing (RT) for intermediate activations and Logit-layer Testing (LT) for the output layer—are calibrated solely on benign data. We then fuse their signals through Recovery-and-Logit Testing (RLT), aligning RT and LT scores via quantile normalization to robustly flag a broad spectrum of attacks, including those that evade one probe in isolation.
@@
The distinguishing features and requirements of our method are summarized in Table 1, where we compare RT and LT against several representative baselines. As shown, our method is self-sufficient (requires no adversarial examples), model-local (requires no external SSL), and low-overhead (no kNN graphs or excessive augmentation), making it suitable for scalable deployment.
@@
  • We propose a novel layer-wise adversarial detection paradigm that exploits partial consistency in internal representations of adversarial examples by focusing on feature transformations. Unlike prior methods, our approach avoids significant overhead by eliminating the need for heavy augmentations, complex structures to maintain reference sets, external large pre-trained models, such as SSL, or large ensemble models.
  • We introduce two practical testing strategies—Recovery Testing and Logit-layer Testing—which identify inconsistencies in intermediate and final-layer representations, respectively, using only the target model’s own feature space.
  • We evaluate our framework on standard benchmarks (CIFAR-10, CIFAR-100, and ImageNet), demonstrating strong detection performance, low computational overhead, and negligible impact on clean inputs, making it practical for real-world deployment.
By focusing on unique internal feature transformations over reference-based comparisons or heavy augmentations, our method avoids extensive retraining and specialized pipelines while delivering a more efficient, robust detector. Next, we review related work, detail our layer-consistency analysis, and present experiments across diverse adversarial threat models.
@@
Ours Baseline Methods Criterion RT LT BEYOND [39] LID [23] Mao [24] Hu [14] DkNN [27] kNN-Def. [9] Self-Sufficient Data? (No Adversarial Data Needed?) Standalone Model? (No Extra Pre-trained Model Dependencies?) No Heavy kNN Retrieval? (No Nearest-Neighbor Search Overhead?) No Excessive Augmentations? (No Many Augmentations?) No Extra Optimization? (No Additional Training Needed?) /
2 Related Work and Background
@@
Recent approaches have explored feature consistency using self-supervised learning (SSL) models. For example, BEYOND [39] detects adversarial inputs by measuring feature stability across augmentations in a large SSL encoder. These methods eliminate the need for adversarial examples and external graph structures but introduce high computational costs due to reliance on large pretrained models, which may be significantly larger than the target classifier itself (e.g., $\times2$ larger parameters). Additionally, they often require access to domain-specific SSL models and retraining or fine-tuning of auxiliary heads.

### v5

@@ Abstract
Deep neural networks (DNNs) are highly susceptible to adversarial examples—subtle, imperceptible perturbations that can lead to incorrect predictions. While detection-based defenses offer a practical alternative to adversarial training, many existing methods depend on external models, complex architectures, or adversarial data, limiting their efficiency and generalizability. We introduce a lightweight, plug-in detection framework that leverages internal layer-wise inconsistencies within the target model itself, requiring only benign data for calibration. Our approach is grounded in the A Few Large Shifts Assumption, which posits that adversarial perturbations induce large, localized violations of layer-wise Lipschitz continuity in a small subset of layers. Building on this, we propose two complementary strategies—Recovery Testing (RT) and Logit-layer Testing (LT)—to empirically measure these violations and expose internal disruptions caused by adversaries. Evaluated on CIFAR-10, CIFAR-100, and ImageNet under both standard and adaptive threat models, our method achieves state-of-the-art detection performance with negligible computational overhead. Furthermore, our system-level analysis provides a practical method for selecting a detection threshold with a formal lower-bound guarantee on accuracy. The code is available here: https://github.com/c0510gy/AFLS-AED.
@@ 1 Introduction
Existing defensive strategies generally fall into three categories: adversarial training, input purification, and AE detection [15, 1]. Adversarial training, re-trains or fine-tunes a model using adversarially perturbed samples to have Adversarially Trained Classifier (ATC) [35, 22], despite its robustness [11, 37], it incurs high computational costs and may compromise clean-data accuracy. Input purification methods [26, 32] attempt to remove adversarial noise through preprocessing (e.g., denoising), yet often fail against adaptive attacks [8].
@@
A more flexible alternative is AE detection, which avoids the costly retraining process of ATCs by instead rejecting suspicious inputs at test time [36, 41]. Detection-based defenses offer distinct advantages, including lower implementation costs, as they do not require exhaustive adversarial training with adversarial data, and tunable robustness, where sensitivity can be adjusted to meet application-specific accuracy–robustness trade-offs. Furthermore, their plug-and-play nature allows them to be integrated with existing models, including ATCs, to further enhance system-level robustness. By architecturally separating the detection mechanism from the classifier, this approach introduces an additional layer of security at the system level. AE detection techniques are often categorized into two types: those that compare inputs to a reference set (e.g., Deep k-Nearest Neighbors, DkNN [29]; Latent Neighborhood Graph, LNG[1]), and those that analyze invariants in the learned representations [17, 7]. Despite these benefits, prominent detection paradigms suffer from practical limitations. Reference-based detectors often require storing adversarial examples or constructing complex neighbor graphs, which is computationally intensive. To address this, recent work [40] has proposed detecting consistency between augmented inputs using pre-trained Self-Supervised Learning (SSL) models. However, this introduces significant overhead from large external models—which can be over 22 times larger than the target classifier—and assumes the availability of high-quality, domain-specific SSL models.
@@
To eliminate the overhead of external models, complex data structures, or heavy augmentations, we develop a self-contained, layer-wise detector that scrutinizes the network’s own representations. This raises a central question: How do adversarial perturbations propagate through a deep neural network? Prior work suggests that such perturbations often leave early-layer features largely intact while inducing sharp deviations in deeper layers. Motivated by these findings, our own empirical analysis confirms this behavior (subsection 3.4). As visualized in Figure 3, we observe that while benign inputs produce relatively flat error profiles across layers, many common adversarial attacks induce large, localized shifts at only a few critical points in the network. Rather than contrasting these deviations with external references, we ask a different core question: Do these internal feature jumps themselves expose adversarial inputs? We formalize this intuition in the A Few Large Shifts Assumption, which posits that adversarial perturbations produce large, localized shifts between a few critical layers. To quantify the observed localized shifts, we introduce the concept of layer-wise Lipschitz continuity, defining these shifts as localized violations of layer-wise Lipschitz continuity. Guided by this principle, our framework is the first designed to empirically measure these violations using a detector calibrated only on benign data. As illustrated in Figure 1, these sparse disruptions leave a detectable footprint across successive layers, which we capture using a lightweight detection framework grounded in layer-wise inconsistency with two complementary probes: Recovery Testing (RT) and Logit-layer Testing (LT). We then fuse their signals through Recovery-and-Logit Testing (RLT), aligning their scores via quantile normalization to robustly flag a broad spectrum of attacks, including those that evade one probe in isolation.
@@
The distinguishing features and requirements of our method are summarized in Table 1, where we compare RT and LT against several representative baselines. As shown, our method is self-sufficient (requires no adversarial examples), model-local (requires no external SSL), and low-overhead (no kNN graphs or excessive augmentation). Furthermore, while prior AE detection works lack the thorough system-level analysis essential for practical deployment, our work addresses this gap by introducing a formal method to select a detection threshold with a guaranteed lower bound on overall system accuracy, as detailed in Appendix H. This enhances the practical reliability of our detector, making it highly suitable for scalable deployment.
@@
  • We propose a novel adversarial detection paradigm that exploits partial consistency in internal feature transformations of adversarial examples. Unlike prior methods that rely on computationally expensive external reference sets or large auxiliary models, our approach eliminates this overhead by measuring inconsistencies across a model’s own layers for a single input.
  • We are the first to formalize these internal disruptions as violations of layer-wise Lipschitz continuity. We introduce two complementary probes, Recovery Testing (RT) and Logit-layer Testing (LT), as practical and efficient methods to empirically quantify these violations using only benign data.
  • We conduct a comprehensive evaluation on standard benchmarks, demonstrating state-of-the-art performance against a wide range of threats, including strong, end-to-end adaptive attacks and generalization to various architectures. Crucially, our system-level analysis yields a formal method for selecting a detection threshold with a guaranteed lower bound on system accuracy, enhancing the method’s practical reliability.
@@
Ours Baseline Methods Criterion RT LT BEYOND LID Mao Hu DkNN kNN-Def. Self-Sufficient Data? (No Adversarial Data Needed?) Standalone Model? (No Extra Pre-trained Model Dependencies?) No Heavy kNN Retrieval? (No Nearest-Neighbor Search Overhead?) No Excessive Augmentations? (No Many Augmentations?) No Extra Optimization? (No Additional Training Needed?) /
2 Methodology
In this section, we introduce our layer-wise adversarial example detection framework (Figure 2), comprising three measures: RT, LT, and their fusion RLT. RT captures inconsistencies in intermediate-layer embeddings, while LT measures logit instability under augmented inputs. At test time, pick a detection measure $m\in\{\mathrm{RT},\mathrm{LT},\mathrm{RLT}\}$ and threshold $\tau_{m}$. We declare $x$ adversarial iff $m(x)>\tau_{m}$, i.e., $\hat{a}(x)\;=\;\mathbb{I}\{\,m(x)>\tau_{m}\,\}\in\{0,1\}.$ Details of the selection of each threshold $\tau_{m}$ are delineated in Appendix H. The algorithm for calculating each measure is shown in Algorithm 1. Theoretical assumptions and proofs of our method are shown in Appendix B and Appendix C, respectively.
2.1 Definition and Notations
2.2 Core Assumption
Assumption 1 (A Few Large Shifts).
This assumption formalizes the intuition that adversarial perturbations cause disproportionately large changes in representation between specific adjacent layers while most other transitions remain relatively stable—a phenomenon that can be quantified based on Lipschitz continuity, which we term layer-wise Lipschitz continuity (See Appendix B.1). Specifically, the condition $\|\mathbf{z}_{k+1}(x^{\prime})-\mathbf{z}_{k+1}(x)\|_{2}\gg\|\mathbf{z}_{k}(x^{\prime})-\mathbf{z}_{k}(x)\|_{2}$ implies a large local Lipschitz constant for the transformation between layers, indicating a region of high instability. The following sections introduce our proposed testing measures, which are designed to empirically identify and quantify these localized violations of layer-wise Lipschitz continuity.
2.3 Recovery Testing Measure
At test time, we define the normalized squared error for each layer $k$ as
Collecting the vector $e(x)=(e_{1}(x),\ldots,e_{L-1}(x))$, we normalize it with softmax to obtain a distribution $\sigma(e(x))$ across layers. The non‐uniformity of this distribution is captured using inverse entropy $\log(L-1)-\mathcal{H}(\sigma(e(x)))$. The final score is computed by taking the average of the raw reconstruction errors and weighting it by this information‐based term:
2.4 Logit-layer Testing Measure
Although RT effectively detects perturbations in intermediate layers $f_{i}$, it cannot be applied to the final logit layer $f_{\text{logit}}$, since there is no subsequent representation from which to reconstruct the logits. To address this—and motivated by the intuition that adversarial examples aim to induce misclassification—we introduce Logit‐layer Testing (LT): we quantify uncertainty at $f_{\text{logit}}$, relative to changes in intermediate features, using data‐driven, low‐cost input augmentations.
Let $\{W^{(g)}\}_{g=1}^{G}$ be a small set (e.g., $1\leq G\leq 6$) of image transformation matrices. Each matrix $W^{(g)}$ is initialized as an identity transformation and is then fine-tuned on benign data using gradient descent. The LT score itself, as defined in Equation 4, serves as the loss function for this process, encouraging the transformations to find augmentations that preserve logit stability for benign inputs, i.e., $\sigma(f(x))\approx\sigma(f(W^{(g)}x))$, while still perturbing intermediate features. This establishes a consistent baseline against which adversarial instability can be measured.
$\mathcal{L}_{LT}=\frac{1}{N}\sum_{n=1}^{N}LT(x_{n}).$ (5)
2.5 Recovery and Logit Testing Combined Measure
Let $\hat{\mathcal{F}}_{RT}$ and $\hat{\mathcal{F}}_{LT}$ be the empirical cumulative distribution functions (CDFs) of RT and LT scores computed over the benign data. Let $\Phi^{-1}$ denote the quantile function (inverse CDF) of the standard normal distribution. At test time, each score is transformed as follows:
$RLT(x)=(RT_{norm}(x))^{2}+(LT_{norm}(x))^{2}.$ (6)
3 Experiments
FGSM PGD AutoAttack Square FGSM PGD CW AutoAttack Square
LNG [1] 98.51 63.14 58.47 94.71 99.88 91.39 89.74 84.03 98.82
Our Approaches
RT 99.93 ±0.005 96.89 ±0.071 99.99 ±0.000 85.38 ±0.344 99.93 ±0.005 96.89 ±0.071 99.99 ±0.002 99.99 ±0.000 85.38 ±0.344
RLT 99.85 ±0.005 99.37 ±0.011 99.99 ±0.000 95.95 ±0.102 99.85 ±0.005 99.37 ±0.011 99.91 ±0.004 99.99 ±0.000 95.95 ±0.102
Method Unseen Seen
FGSM PGD FGSM PGD CW
kNN [10] 51.63 ±0.04 51.14 ±0.039 51.63 ±0.04 51.14 ±0.039 50.73 ±0.04
LNG [1] 96.85 89.61 99.53 98.42 86.05
Our Approaches
RT 94.31 ±0.457 99.99 ±0.000 94.31 ±0.457 99.99 ±0.000 92.18 ±0.135
LT 96.18 ±0.028 97.89 ±0.021 96.18 ±0.028 97.89 ±0.021 94.06 ±0.215
RLT 97.60 ±0.048 99.99 ±0.000 97.60 ±0.048 99.99 ±0.000 91.19 ±0.022
Defense $\bm{L_{\infty}=0.01}$ $\bm{L_{\infty}=8/255}$
RA@FPR5% RA@FPR50% RA@FPR5% RA@FPR50%
Ours (RLT) 75.40 99.58 33.70 80.77
BEYOND [40] 88.38 98.81 13.80 48.20
Trapdoor [31] 0.00 7.00 0.00 8.00
DLA [33] 62.60 83.70 0.00 28.20
SID [34] 6.90 23.40 0.00 1.60
SPAM [21] 1.20 46.00 0.00 38.00
Table 5: RA (%) scores of end-to-end attack with PGD under perturbation $\epsilon=8/255$ (ResNet110, CIFAR10). The results are the mean and standard deviation of 5 runs.
$\bm{\lambda}$ @FPR5% @FPR10% @FPR15% @FPR20% @FPR25% @FPR30% @FPR35% @FPR40% @FPR45% @FPR50%
1.00 $20.72_{\pm 8.04}$ $25.96_{\pm 6.88}$ $32.05_{\pm 5.72}$ $38.23_{\pm 5.11}$ $44.48_{\pm 4.52}$ $0.15_{\pm 4.13} 55$$5.15_{\pm 3.56}$ $60.09_{\pm 3.10}$ $64.54_{\pm 2.42}$ $68.53_{\pm 2.21}$
0.50 $22.64_{\pm 4.46}$ $30.45_{\pm 3.25}$ $37.96_{\pm 2.08}$ $44.78_{\pm 1.81}$ $51.19_{\pm 1.30}$ $56.41_{\pm 0.98}$ $61.21_{\pm 0.96}$ $65.44_{\pm 0.96}$ $69.35_{\pm 0.83}$ $73.00_{\pm 0.92}$
0.25 $23.39_{\pm 1.76}$ $32.92_{\pm 0.49}$ $40.88_{\pm 1.04}$ $48.04_{\pm 1.13}$ $54.32_{\pm 1.58}$ $59.37_{\pm 1.44}$ $63.90_{\pm 1.82}$ $67.98_{\pm 1.82}$ $71.51_{\pm 1.89}$ $74.83_{\pm 1.85}$
Mao [26] 18.97
BEYOND [40] 19.45
Table 6: RA (%) scores of end-to-end attack with PGD under perturbation $\epsilon=8/255$ (ResNet110, CIFAR10) when using an ATC. The results are the mean and standard deviation of 5 runs.
$\bm{\lambda}$ @FPR5% @FPR10% @FPR15% @FPR20% @FPR25% @FPR30% @FPR35% @FPR40% @FPR45% @FPR50%
1.00 $93.32_{\pm 0.13}$ $93.38_{\pm 0.14}$ $93.41_{\pm 0.11}$ $93.43_{\pm 0.09}$ $93.97_{\pm 0.12}$ $94.49_{\pm 0.15}$ $94.99_{\pm 0.14}$ $95.50_{\pm 0.13}$ $95.93_{\pm 0.14}$ $96.32_{\pm 0.11}$
0.50 $93.53_{\pm 0.15}$ $94.17_{\pm 0.12}$ $94.27_{\pm 0.11}$ $94.33_{\pm 0.09}$ $94.86_{\pm 0.13}$ $95.41_{\pm 0.09}$ $95.88_{\pm 0.09}$ $96.27_{\pm 0.11}$ $96.65_{\pm 0.14}$ $97.00_{\pm 0.15}$
0.25 $93.57_{\pm 0.14}$ $94.46_{\pm 0.16}$ $94.60_{\pm 0.13}$ $94.85_{\pm 0.07}$ $95.40_{\pm 0.05}$ $95.87_{\pm 0.08}$ $96.32_{\pm 0.14}$ $96.71_{\pm 0.12}$ $97.04_{\pm 0.12}$ $97.30_{\pm 0.14}$
Mao [26] 75.09
BEYOND [40] 93.20
Table 7: Evaluation of RA (%) on end-to-end gradient-free SimBA attack with 1000 steps under perturbation $\epsilon=8/255$ (ResNet110, CIFAR10).
Defense RA@FPR5% RA@FPR50%
None 5.54
Ours (RLT) 97.62 97.85
[…]
Mao [26] 5.25 38.12 22.02
LNG [1] 0.286 8.33 4.81
BEYOND [40] 0.715 20.62 11.91
Ours (RLT) 0.491 2.59 1.49
3.1 Experimental Setup
Datasets. We evaluate our method on both CIFAR-10 and ImageNet datasets. CIFAR-10 consists of 60,00060{,}000 $32\times 32$ images across 10 classes, with standard training (50k) and test (10k) splits. For ImageNet, we use the official training and validation sets, resizing all images to $256\times 256$ and applying standard normalization. CIFAR-10 is used to validate the generality and scalability of our detection approach in a lower-resolution setting, while ImageNet serves as a large-scale benchmark.
3.2 Detection Performance under Standard Attacks
3.3 Detection Performance under Adaptive Attacks
To avoid gradient obfuscation and guarantee end-to-end gradient flow through our detection framework, we apply the Backward Pass Differentiable Approximation (BPDA) [4] to components that may otherwise block gradients, such as the quantile module.
We first evaluate robustness under the Orthogonal-PGD adaptive attack following BEYOND [40], where the adversary has full knowledge of both the classifier and our detection mechanism and explicitly optimizes to induce misclassification while suppressing detection. Table 4 reports robust accuracy (RA) at two $L_{\infty}$ budgets ($0.01$ and $8/255$) and operating points (FPR=5% and FPR=50%). While BEYOND benefits from additional SSL components, our method attains competitive or superior RA using only model-internal signals and benign calibration, without adversarial data or external encoders. This demonstrates that even when gradients are orthogonalized, the detection terms themselves impose inherently conflicting objectives, thereby constraining the attacker’s effectiveness.
We further evaluate fully end-to-end PGD attacks on the fused RLT score to validate the analysis on objective confliction in Appendix B.5.2, using the untargeted loss $\mathcal{L}_{\mathrm{cls}}(x+\delta,y)+\lambda\cdot\mathrm{RLT}(x+\delta)$. Table 5 shows that, although RA decreases under this strong adaptive threat, our method consistently outperforms baselines’ best performance report at $\mathrm{FPR}{=}5\%$, such as Mao et al. and BEYOND across all tested $\lambda$ and FPRs. When combined with an ATC (Table 6), the system remains highly resilient, achieving over $93.3\%$ RA at $\mathrm{FPR}{=}5\%$, again surpassing the adaptive-attack results of Mao et al. and BEYOND at $\mathrm{FPR}{=}5\%$. These findings underscore that even under fully end-to-end optimization with varying attack strengths $\lambda$, robustness is maintained through gradient conflicts induced by our fused detection terms.
To rule out gradient masking completely, we evaluate a query-only, gradient-free attack (SimBA) [14] with 1,000 queries. Table 7 shows that although SimBA reduces the undefended classifier’s RA to 5.54%, our RLT detector restores RA to 97.62% at 5% FPR. This strong performance against a computationally intensive black-box attack confirms that our defense does not rely on obfuscated gradients and remains robust beyond first-order threat models.
3.4 Empirical Evaluation of the Proposed Assumption
3.5 Implementation Costs
We evaluate the computational efficiency of our RLT method relative to existing baselines under the same configuration as in Table 2. As shown in Table 8, RLT introduces minimal floating-point operations (FLOPs), parameters, and model size overhead, offering a lightweight, plug-in solution without retraining or external models. It is significantly more efficient than SSL-based (Mao et al.[26]) and graph-based (LNG[1]) detectors, making it well-suited for real-time and resource-constrained deployments. A detailed analysis of this detection overhead across various architectures is provided in Appendix G.
4 Conclusion
In this work, we introduced a novel framework for adversarial example detection that leverages layer-wise inconsistencies within deep neural networks. Motivated by our proposed A Few Large Shifts Assumption—which posits that adversarial perturbations cause large, localized violations of layer-wise Lipschitz continuity—we developed two complementary detection strategies: RT and LT. Extensive evaluations on CIFAR-10, CIFAR-100, and ImageNet demonstrate that our combined approach, RLT, achieves state-of-the-art detection performance against a wide range of threats, including strong adaptive attacks. Our method operates efficiently without relying on external models or extensive augmentation, and our system-level analysis provides a practical method for selecting a detection threshold with a formal lower-bound guarantee on accuracy, highlighting its suitability for real-world deployment.
Limitations and Future Work. Despite these results, the framework’s effectiveness relies on an assumption grounded in empirical observations of current attacks and thus cannot guarantee defense against all possible future attacks. Key directions for future work therefore include deepening the theoretical understanding of layer-wise Lipschitz continuity to determine if the "A Few Large Shifts" principle is a fundamental property of adversarial examples, and developing more sophisticated methods for analyzing the challenging logit space. This exploration could lead to provably robust defenses.
Acknowledgements
Appendix A Related Work and Background
@@
Recent approaches have explored feature consistency using self-supervised learning (SSL) models. For example, BEYOND [40] detects adversarial inputs by measuring feature stability across augmentations in a large SSL encoder. These methods eliminate the need for adversarial examples and external graph structures but introduce high computational costs due to reliance on large pretrained models, which may be significantly larger than the target classifier itself (e.g., $22\times$ larger parameters). Additionally, they often require access to domain-specific SSL models and retraining or fine-tuning of auxiliary heads.